OpenStudy (anonymous):

The sum of the ages of a father and son is 56. Four years ago the father was 3 times as old as the son. Find thepresent age of each.

OpenStudy (lgbasallote):

let x= father's age 56 - x = son's age x-4 = 3(56-x -4) x - 4 = 3(52 - x) can you take over from here?